For the year ended December 31, 2020, the job cost sheets of Cinta Company contained the following data. Job Number 7640 Expl• Your answer is incorrect. Determine the gross profit to be reported for 2020. Gross profit $
Answer #1

Solution:

Sales = $583,000

Cost of Jobs sold 7638, 7639, 7641 = 95700 + 101200 + (55660+163460) = $416,020

Manufacturing Overhead overapplied = Applied overhead - Actual Overhead

= (47520+63360+72600) - (132000+15400+19800+8800) = $7,480

Gross Profit = Sales - Cost of Jobs sold + Manufacturing Overhead overapplied

= $583000- 416020+ 7480 = $174,460